OpenStudy (anonymous):

6x+3y=36 find the y and x intercepts

OpenStudy (anonymous):

x-intercept is a point on the graph where y is zero. y-intercept is a point on the graph where x is zero. y = 0 6x+3y=36 6x+3*0=36 6x=36 x = 6 (6,0) x = 0 6x+3y=36 6*0+3y=36 3y = 36 y = 12 (0,12)

OpenStudy (anonymous):

To add to what @surdawi is saying, To find a y - intercept, set x = 0 and solve. (functions will ALWAYS just have 1) x -intercept, set y = - 0 and solve. (functions usually will have the number of intercepts as the highest degree in the function)